OpenStudy (anonymous):

MULTIPLYING RADICALS: sqrtt10 * sqrt 8

OpenStudy (anonymous):

\[\LARGE \sqrt{10} * \sqrt{8}\]

OpenStudy (anonymous):

You simply multiply them after simplification. \[\sqrt{10}*\sqrt{8}=\sqrt{10}*(2\sqrt{2})=2\sqrt{20} = 2\sqrt{2*2*5}=2*2\sqrt{5}=4\sqrt{5}\]

OpenStudy (anonymous):

thank you :)